Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by kartikshah » Fri Jul 27, 2012 6:07 pm
What is the probability that a 3-digit positive integer picked at random will have one or more "7" in its digits?

A. 271/900

B. 27/100

C. 7/25

D. 1/9

E. 1/10

Apart from providing solution to the above question, could someone guide me to a good resource on probability, permutations and combinations? I'm going to appear for my exam early next week but I still don't feel confident about this topic!!
Join the discussion
Source: — Problem Solving |

by eagleeye » Fri Jul 27, 2012 6:16 pm
kartikshah wrote:What is the probability that a 3-digit positive integer picked at random will have one or more "7" in its digits?

A. 271/900

B. 27/100

C. 7/25

D. 1/9

E. 1/10

Apart from providing solution to the above question, could someone guide me to a good resource on probability, permutations and combinations? I'm going to appear for my exam early next week but I still don't feel confident about this topic!!
Total numbers from 100 to 999 = 999-100+1 = 900
first digit cant be 0.
Digits with 7 missing = (8 options for 1st digit)*(9 options for 2nd)*(9 options for 3rd)
= 8*9*9

Required probability = 1 - 8*9*9/(900) = 1- 18/25 = 7/25.

Unfortunately I don't have a specified resource, but you should check out Brent's GMATPrepNow. He posts really good probability solutions on the forum.
Join the discussion

by kartikshah » Fri Jul 27, 2012 6:22 pm
I didn't get why you subtracted 648/900 from 1
When we say
d1 = 8 ways to choose; doesn't that include 7
d2 = 9 ways to choose; doesn't that include 7
d3 = 9 ways to choose; doesn't that include 7

?
Please explain.
Join the discussion

by eagleeye » Fri Jul 27, 2012 6:31 pm
kartikshah wrote:I didn't get why you subtracted 648/900 from 1
When we say
d1 = 8 ways to choose; doesn't that include 7
d2 = 9 ways to choose; doesn't that include 7
d3 = 9 ways to choose; doesn't that include 7

?
Please explain.
No. None of them include 7. Total number of possible digits for d2 and d3 = 10 (0,1,2,3,4,5,6,7,8,9). If I de-select 7, I get 9 digits as you saw.
Does this work?
Join the discussion

by kartikshah » Fri Jul 27, 2012 6:32 pm
Yea, you're right!
I had not counted zero as a possible digit for d2 and d3!!
Join the discussion